### v1

# Balancing Performance and Efficiency: A Multimodal Large Language Model Pruning Method based Image Text Interaction

Abstract
Recently, multimodal large language models (MM-LLMs) have achieved great success in many multimodal tasks, but their high computational costs limit their further promotion and application. In the MM-LLMs framework, the main computational consumption step is the processing of concatenated text and visual tokens at the LLM layer. The length of the input token for LLM directly affects the overall training and inference efficiency. In response to this issue, we further studied the visual tokens of MM-LLMs. We found that the similarity between visual and CLS tokens in the visual encoder follows a long-tail distribution. In other words, only a few visual tokens are highly similar to CLS tokens. Therefor, we designed a dynamic pruning algorithm to address this issue. Firstly, for different input samples, we search for the inflection point of their visual CLS token similarity curve and use it as the corresponding segmentation point to trim the visual markers. This process mainly reduces the output of the visual encoder to accelerate the model. Then, in the LLM layer, the concatenated visual text tokens are pruned for the second time. During this process, due to the interaction between visual and textual features, visual and textual tokens with low text correlation are further filtered, achieving a balance between efficiency and performance. The results on multiple datasets show that our proposed method can achieve performance that competes with the original performance when using an average of of the original token quantity. Our source code will be made publicly available following acceptance.
Introduction
Benefiting from the training on massive text data and the continuous development of deep neural network structures, as well as the methods of pre-training and fine-tuning, large language models (LLMs) (Brown et al. 2020; Ouyang et al. 2022; Jiang et al. 2023a; Touvron et al. 2023) have become the forefront of research in the field of artificial intelligence. Despite the remarkable achievements of deep learning methods in various tasks in recent years, the emergence of LLMs has sparked a revolutionary breakthrough in the field of artificial intelligence.
Utilizing large language models as the core component, multimodal large language models (MM-LLMs) (Liu et al. 2023a, 2024a; OpenAI 2023; Team et al. 2023) take both visual and textual information as input to address multimodal tasks. Specifically, MM-LLMs employ a visual encoder, such as Vision Transformer (ViT) (Dosovitskiy et al. 2021) and CLIP (Radford et al. 2021), and mapping layers, such as Multi-Layer Perceptions (MLP), to extract visual features and transform them into textual feature space. They then leverage the powerful language generation, zero-shot learning, and contextual learning capabilities of LLMs to process concatenated visual and textual features. MM-LLMs demonstrate performance surpassing traditional models across various multimodal tasks, showcasing remarkable emergent capabilities.
Due to the enormous number of parameters in MM-LLMs, both training and inference require substantial computational resources, with LLMs predominantly driving the computational costs due to their larger parameter count compared to the visual encoder. Furthermore, since LLM mostly employs transformer architectures, their computational cost often quadratically increases with input length. Therefore, the efficiency of MM-LLMs is significantly influenced by the number of visual tokens generated by the visual encoder. In addressing the issue of enhancing LLM efficiency, two widely used programmes are currently employed. The first program involves replacing high-parameter LLM with smaller-parameter LLM. By directly reducing the model parameters, significant reductions in computational requirements can be achieved, thereby speeding up both training and inference. For instance, Chu et al (Chu et al. 2023). proposed a mobile device MM-LLM by replacing the LLM with and parameters model. To achieve local deployment, Yuan et al (Yuan, Li, and Sun 2023). proposed the TinyGPT-V, based on small backbones with a parameter. Although replacing the original large parameter LLM with lightweight LLM can improve the computational efficiency and local deployment of the model, this method will reduce the inference ability of the model, resulting in a significant decrease in performance.
The second program involves reducing the output of the visual encoder to decrease the input length of LLM, effectively enhancing the overall training and inference speeds of the model. Recently, there have been related works that use this to achieve efficient inference of MM-LLMs. For example, Wang et al (Wang et al. 2024). designed a lightweight module for identifying and removing redundant tokens and attention heads in each layer. By adding this module to the original model, the training and inference process of the model can be accelerated. Shi et al (Shi et al. 2024). presents Cross-Guided Ensemble of Tokens (CrossGET), a versatile acceleration framework designed for vision-language Transformers. By dynamically merging tokens during inference, CrossGET effectively lowers computational expenses without compromising performance quality. Shang et al (Shang et al. 2024). proposed an adaptive token pruning strategy based on the LLaVA framework, which reduces the output of visual encoders by clustering visual tokens, thereby improving the computational efficiency of LLaVA.
Inspired by the second method, we have designed a dynamic language-text interactive pruning method. Specifically, we explored the distribution of visual tokens. As shown in Figure 1, when arranging the similarity between the CLS and the visual token in descending order, we observed a long-tail distribution in the similarity curve. This implies that there is a significant amount of redundancy in visual tokens, with only a small portion of visual tokens highly correlated with the input image. Therefore, we have devised an objective function that compresses visual tokens by aiming to find a segmentation point between the “head” and “tail” of the long-tail distribution. By optimizing this objective function, we can reduce the output of the visual encoder. Subsequently, to further accelerate the model, we have adopted a method similar to Zhang et al (Zhang et al. 2024), where pruning is performed by calculating the importance between mapped visual tokens and textual tokens. During this process, there is an interaction between visual and textual information, and the pruning process can be seen as extracting tokens with high visual-textual relevance. This ultimately leads to a further reduction in the input sequence length of the LLM, enhancing the inference efficiency of MM-LLMs. At the same time, due to the different distribution of visual tokens for different tasks and instances, we have added a dynamic pruning rate setting to dynamically process each sample/task during the pruning process, ensuring that the model can achieve a balance between efficiency and performance on different data.
Our main contributions are summarized as follows:
We have designed a text visual interaction pruning algorithm. Firstly, we reduce the output of the visual encoder based on our designed objective function. Then, we calculate the importance of the visual/text tokens before inputting them to LLM and extract highly correlated visual and text tokens based on this. This achieves competitive performance with a compression ratio of no less than 8x.
Based on the different input instances, we have proposed the dynamic pruning rate to ensure dynamic pruning for each sample/task, thus ensuring the generalization of the algorithm.
Our method is simple and efficient to implement, not only can it be directly incorporated into the model for training-free inference, but it can also participate in training with the model during the fine-tuning phase.
Related Work
Visual Token Pruning Method
In the MM-LLMs framework, visual encoders commonly leverage the ViT framework. Presently, numerous researchers are dedicated to enhancing the efficiency of ViT models, with a focus on token compression emerging as a key area of study. Token compression strategies encompass both token pruning and token merging techniques.
The former evaluates the importance of different tokens based on defined criteria, retaining significant tokens while discarding insignificant ones. For instance, Rao et al (Rao et al. 2021). introduced a dynamic pruning technique that gradually and adaptively prunes redundant tokens at each model layer by considering the visual attention sparsity. This method estimates the importance scores of individual tokens using current features. Going beyond this, Kong (Kong et al. 2022) and Xu (Xu et al. 2023) proposed that insignificant tokens should not be merely discarded but instead incorporated or subject to additional modifications to prevent the permanent loss of image data resulting from improper pruning.
The latter involves grouping similar tokens together, discarding unimportant background tokens, and achieving efficient token compression by merging foreground tokens. Chen et al (Chen et al. 2023). linked the loss function with the compression rate to autonomously determine varied token compression rates across different layers, employing a combination of pruning and merging techniques simultaneously. Long et al (Long et al. 2023). ensured the merging process’s reliability by factoring in token importance and diversity for pruning, subsequently merging akin tokens. Similarly, Lee et al (Lee, Choi, and Kim 2024). underscored the significance of considering diverse token relationships during merging processes. They devised multiple criteria for the gradual fusion of tokens, striking an optimal balance between speed and precision.
Efficient Large Language Model
Large language models have garnered significant attention due to their outstanding performance across various tasks. However, the substantial computational and memory requirements for inference with large language models pose challenges for their deployment in resource-constrained environments. Researchers have been diligently working on developing technologies aimed at enhancing the efficiency of inference with large models.
Methods to enhance the efficiency of large models can be broadly categorized into three levels of optimizations: data-level, model-level, and system-level (Zhou et al. 2024). Data-level optimization improves efficiency by optimizing input prompts, without altering the original model, thereby avoiding high model training costs. For example, Jiang et al (Jiang et al. 2023b). divided the prompt into three parts, and then used additional models to calculate the perplexity of each part separately. Based on the perplexity, prompts were removed to accelerate the model. On this basis, Jiang et al (Jiang et al. 2023c). proposed using question text to compress the document section, further improving the computational efficiency of the model.
Model-level optimization is typically carried out during the model inference process, by designing effective model structures or compressing pretrained models to improve efficiency. Usually, it requires pretraining or fine-tuning to preserve or restore model functionality. Such as Zhang et al (Zhang et al. 2024). found that only a small number of tokens made outstanding contributions during the attention process. Based on this, they designed a dynamic KV cache eviction policy that balances recent and important tokens. Li et al (Li et al. 2024). discover that each attention head in the model consistently focuses on specific prompt attention features during generation. Drawing on this insight, they proposed the SnapKV, which automatically compresses KV caches by selecting clustered important KV positions for each attention head.
System-level optimization involves optimizing the inference engine or service system. Optimization of the inference engine does not require model training, and optimization of the service system is non-destructive to model performance.
Building upon the works above, we have devised a multimodal dynamic token pruning method that integrates text and visual interactions. Initially, we observed a long-tail distribution of similarity between the visual and the CLS token generated by the visual encoder. This suggests that not all tokens contain valuable information for the input image. Hence, we developed a dynamic segmentation function to locate the split point between the “head” and “tail” of this long-tail distribution, extracting important tokens in the visual modality for pruning, ensuring efficiency, and maintaining high performance. Subsequently, at the LLM layer, we computed the importance of visual tokens aligned with the text modality and text tokens mutually. Based on this evaluation, we pruned the inputs to the LLM layer once more, striking a balance between efficiency and performance. Our approach dynamically processes each input instance, showcasing strong generalization capabilities.
Method
Overview
In order to improve computational efficiency while ensuring performance availability, we have designed a visual-text interactive dynamic pruning method as shown in Figure 2. This method consists of the following three steps:
Visual Token Pruning Firstly, calculate the similarity between CLS and visual tokens and sort the visual tokens according to their similarity. Then, maximize the dynamic segmentation function to extract tokens with high CLS-visual similarity and discard other tokens to complete the pruning of the visual tokens.
Token mapping alignment and concatenation The pruned visual token is aligned to the feature space of the textual token through a mapping layer. And concatenate them with the text tokens.
Token pruning based on visual-text interaction Compute the similarity between each token and its preceding token, using this as the importance score for each token. Sort based on this importance and remove the lowest-scoring N tokens to reduce the input length at the LLM layer. During the calculation of importance scores, for each current token, the similarity with all preceding tokens is considered. Therefore, for every text token, similarity with all visual tokens is computed, biasing towards retaining tokens with high visual-text correlation.
After completing the above steps, send the token to LLM to obtain the final output.
Dynamic Visual Token Pruning Algorithm
Inspired by the works of Zhang (Zhang et al. 2024) and Li (Li et al. 2024), we conducted a study on the similarity between the CLS and visual tokens generated by the visual encoder. Our research revealed the sparsity of similarity between the CLS and visual tokens, indicating that only a few tokens exhibit high similarity. Therefore, we arranged the similarity scores of CLS-visual token pairs in descending order. As shown in Figure 3, the curve demonstrates a long-tail distribution, suggesting that it is possible to retain only the “head” portion containing tokens with high similarity, while discarding the “tail” portion with low similarity to reduce the output length of the visual encoder and improve overall computational efficiency.
To segment the long-tail distribution curve effectively, firstly, calculate the similarity between CLS and visual tokens, and the similarity formula is as follows:
Then, arrange the similarity in descending order, the CLS-visual similarity set can be obtained as and is the number of tokens.
The definition of the splitting objective function is as follows:
wherein and .
To obtain the optimal segmentation point (i.e. the segmentation position of the ”head” and ”tail”), it is necessary to solve for the solution that maximizes the Equation 2. The above can be defined as:
The meaning of this Equation 2 and 3 is as follows:
For each possible segmentation point (from to ), we calculate .
The molecule of represents is the number of data points to the right of the segmentation point (including the segmentation point) and is the difference between the maximum value and the segmentation point value.
The denominator is a constant used for normalization.
We search for the value of that maximizes this ratio, which is the optimal segmentation point.
To prevent gradient collapse during model training, is multiplied by a smoothing coefficient in practical implementation. The overall steps are shown in Figure 4. After the operation above, the pruning of the visual modality is completed. It is important to note that this process is dynamic for each input instance. This means that for different inputs, we will individually solve the objective function based on the distribution of their visual tokens. This dynamic approach for each sample balances efficiency and effectiveness by providing targeted processing for each instance.
Visual-Text Interactive Token Pruning Algorithm
After pruning the output of the visual encoder, inspired by (Zhang et al. 2024), we adopt a similar approach to further enhance the computational efficiency. We will perform a second round of token pruning before feeding into LLM.
Firstly, the pruned output of the visual encoder will be passed through a projection layer to map it to the feature space of text tokens. Next, as shown in Figure 5, the mapped visual tokens will be concatenated with text tokens. Subsequently, the similarity between each token and all preceding tokens will be calculated, where the importance score of each token is the sum of its similarity scores with other tokens. Finally, based on the importance scores, tokens with low importance scores will be removed. The specific drop steps are as follows:
Set the recent tokens that need to be saved and the remaining tokens that need to be retained, where the number of is and the number of is .
Keep the last tokens without dropping.
When the current number of tokens does not exceed , continuously accumulate the score of .
When the current number of tokens , start dropping .
Continue to accumulate . Once the number of tokens , continue the above process.
During this process, the similarity between visual tokens will be calculated once again to complete the second round of pruning in the visual modality. Additionally, each text token will undergo similarity calculations with preceding visual tokens. This implies that visual tokens and text tokens will interactively compute each other’s importance during this process, guiding the model to retain tokens with high similarity between visual and text modalities, thereby further enhancing the computational efficiency of the model.
Experiments
Dataset
Implementation Details
All experiments were conducted in the PyTorch framework on four NVIDIA 4090 24G GPUs. We utilized the LLaVA1.5-7B as our baseline. The visual encoder is CLIP with ViT as the backbone, where the input image size is 336x336 and the patch size is 14x14. Set hyperparameter to 0.075 in dynamic visual token pruning algorithm, while and are set to and in visual-text interactive token pruning algorithm. It should be noted that our visual pruning method is added after the third layer of the CLIP visual encoder. The LoRA parameters are based on the official parameters provided by LLaVA1.5-7B.
Ablation Study
Efficiency for Dynamic Visual Token Pruning
To verify the effectiveness of our proposed dynamic pruning method, we conducted comparative experiments on SQA, TextVQA, POPE, and MME datasets with the method of taking the similarity.
| Method | POPE | TextVQA | MME |
| Topk | 86.5 (218) | 54.97 (272) | 1417.5 (222) |
| Ours | 86.5 (195) | 55.28 (251) | 1426.5 (200) |
| Method | POPE | TextVQA | MME |
| Visual | 86.5 (196) | 55.28 (254) | 1426.5 (201) |
| Visual + Text | 86.5 (196) | 55.26 (254) | 1430.5 (201) |
| Method | ScienceQA | POPE | MMBench | VQAv2 | TextVQA | MME |
| Training-Free | 69.75 | 86.50 | 61.30 | 75.50 | 55.26 | 1430.50 |
| Fine-tuning | 70.15 | 85.50 | 66.00 | 76.80 | 56.61 | 1488.50 |
The results are shown in Table 1. The number in parentheses represents the number of visual tokens used, and the original total number of visual tokens is . On the TextVQA and MME datasets, the dynamic pruning method not only outperforms the fixed TopK method in terms of performance but also excels in efficiency. Specifically, on the POPE dataset ours achieves the same performance with fewer tokens compared to the fixed TopK method. These results align with our findings that only a small portion of visual tokens are effective, suggesting redundancy within visual tokens. This also underscores the effectiveness of the proposed dynamic visual token pruning method.
Efficiency for Visual-Text Interactive Token Pruning
Similarly, to demonstrate the effectiveness of our proposed visual-text interactive token pruning method, we conducted comparative experiments with only visual dynamic pruning. The experimental results are shown in Table 2. The number in parentheses represents the total number of tokens used (i.e. visual + textual tokens).
| Method | MM-Vet | LLaVA-Bench-in-the-Wild |
| LLaVA | 30.5 (714) | 67.8 (772) |
| Visual | 32.2 (266) | 64.9 (355) |
| Visual + Text | 31.5 (191) | 66.2 (192) |
| Method | ScienceQA | POPE | MMBench | VQAv2 | TextVQA | MME |
| BLIP-2 | 61.00 | 85.30 | - | 41.00 | 42.50 | 1293.80 |
| InstrucBILP | 60.50 | - | 36.00 | - | 50.10 | - |
| InstrucBILP | 63.10 | 78.90 | - | - | 50.70 | 1212.80 |
| Shikra | - | - | 58.80 | 77.40 | - | - |
| IDEFICS-9B | - | - | 48.20 | 50.90 | 25.90 | - |
| IDEFICS-80B | - | - | 54.50 | 60.00 | 30.90 | - |
| Qwen-VL | 67.10 | - | 38.20 | 78.80 | 63.80 | - |
| LLaVA-1.5 | 68.40 | 86.40 | 66.10 | 79.10 | 58.20 | 1476.90 |
| Training-Free Method | ||||||
| LLaVA-PruMerge | 68.52 | 70.70 | 56.78 | 65.90 | 53.51 | 1191.50 |
| Ours | 69.75 | 86.50 | 61.30 | 75.50 | 55.26 | 1430.50 |
| Fine-tuning Method | ||||||
| LLaVA-PruMerge | 68.50 | 76.30 | 60.90 | 72.00 | 56.00 | 1350.30 |
| LLaVA-PruMerge+ | 68.30 | 84.00 | 64.90 | 76.80 | 57.10 | 1462.40 |
| CrossGET | 66.70 | 83.90 | 64.70 | 77.30 | 54.90 | 1510.20 |
| Ours | 70.15 | 85.50 | 66.00 | 76.80 | 56.61 | 1488.50 |
The experimental results indicate that there is minimal performance difference between the two settings. Upon analyzing the datasets, it was found that this is due to all answers in the dataset being single-word answers, which results in a small stride for visual-text pruning. However, in practical, answers in MM-LLMs are predominantly long sentences. To further validate our method, we conducted ablation experiments on two datasets with long answers: MM-Vet (Yu et al. 2023) and LLaVA-Bench-in-the-Wild (Liu et al. 2024b). The result is shown in Table 4, demonstrate that visual-text interactive pruning achieves highly competitive performance while significantly reducing the total number of tokens. Moreover, on the LLaVA-Bench-in-the-Wild dataset, our method surpasses approaches that solely prune visual tokens while using fewer tokens. In conclusion, our proposed method can further reduce computational costs while ensuring performance levels.
Fine-tuning VS Training-Free
Due to the simplicity of our method and its lack of additional parameters, it can be seamlessly integrated into multimodal models at any stage, including both the inference and fine-tuning phases. We compared the performance of the method in two scenarios: the inference stage without training and the fine-tuning. The experimental results are shown in Table 3. The experimental results show that fine-tuning the model together with our method can further improve the performance. This aligns with common sense and intuition, as the fine-tuning process aids the LLM in learning to understand the pruned tokens, thereby boosting the model’s performance even more with fewer tokens.
Main Result
To further validate the effectiveness of our method, we conducted comparative experiments with other MM-LLMs and existing MM-LLM token pruning methods. The results are shown in Table 5.
From the experimental results, it is evident that under the training-freed setting, our proposed method achieves the best performance. In terms of efficiency, our method averages 127 visual tokens across the six datasets. Under the fine-tuning setting, our method reaches optimal performance on the ScienceQA, POPE, and MMBench datasets, and competitive performance on the VQAv2, TextVQA, and MME datasets. Compared to LLaVA-PruMerge+, our method offers higher computational efficiency (visual tokens: 127 vs 144), and since we also prune text tokens, we hold an advantage in the total token count. In comparison to CrossGET, our method also demonstrates efficiency advantages. Due to fine-tuning enabling the model to adapt to pruned visual tokens, it may achieve better performance on tasks like ScienceQA and MME datasets. Overall, our proposed method strikes a balance between performance and efficiency, showcasing strong competitiveness and potential.
To further evaluate the computational efficiency of our proposed method, we employed open-source simulation software (Yuan et al. 2024) to simulate the computational efficiency on an NVIDIA A100 GPU. During the reverse simulation calculations, we assumed text tokens as input and the average number of visual tokens used by each method across the six datasets as the input visual token count. The experimental results are shown in Table 6.
| Method |
Quantization |
| LLaVA1.5 | Vicuna-7B | FP16 | 8.5 | 30.3 | 22.2 | 4.1 | ||||||||||
| LLaVA-PruMerge+ | Vicuna-7B | FP16 | 2.7 | 9.9 | 15.4 | 0.9 | ||||||||||
| Ours | Vicuna-7B | FP16 | 2.3 | 9.7 | 13.2 | 0.8 | ||||||||||
| LLaVA1.5 | Vicuna-7B | INT8 | 4.3 | 15.2 | 11.1 | 2.0 | ||||||||||
| LLaVA-PruMerge+ | Vicuna-7B | INT8 | 1.4 | 5.0 | 7.7 | 0.5 | ||||||||||
| Ours | Vicuna-7B | INT8 | 1.2 | 4.9 | 6.6 | 0.4 | ||||||||||
| LLaVA1.5 | Vicuna-7B | INT4 | 2.1 | 14.2 | 5.6 | 1.0 | ||||||||||
| LLaVA-PruMerge+ | Vicuna-7B | INT4 | 0.7 | 2.5 | 3.9 | 0.2 | ||||||||||
| Ours | Vicuna-7B | INT4 | 0.6 | 2.4 | 3.3 | 0.2 |
Compared to baseline methods, our approach demonstrates significant advantages in computational efficiency. Additionally, when compared to other pruning methods with similar performance levels, our method also exhibits slightly higher computational efficiency. This means that our proposed method can effectively enhance computational efficiency while maintaining performance levels.
Visualization
To visually illustrate our method, we conducted visualization experiments, as shown in Figure 6. The grey area represents the regions corresponding to pruned visual tokens, while the areas with normal pixels represent the regions corresponding to retained visual tokens.
The visualization results demonstrate that visual tokens corresponding to background regions have been mostly discarded, while the retained visual tokens are focused on the main objects in the image (such as boats, people, signs, ice cream, etc.). This indicates that our proposed method can guide the model to concentrate on key areas of the image, effectively enhancing computational efficiency.
Conclusion
In this paper, we propose a multimodal large model dynamic pruning method for text-visual interaction. Firstly, we analyze the similarity of CLS visual tokens in the visual encoder, which exhibits a long tail distribution. Therefore, we design an objective function to segment the head and tail in the long tail distribution to extract important visual tokens. Subsequently, the similarity between all tokens is calculated at the LLM layer, and the importance score of each token is obtained by summing up the similarity between the current token and other tokens. Tokens with high importance scores are retained, while tokens with low importance scores are discarded to complete interactive pruning.
The experimental results on multiple datasets showed that our method achieved the best performance on most datasets and maintained a balance between performance and computational efficiency. This provides ideas and methods for future research on efficient multimodal large models
Limitation
In terms of computational efficiency, there is still room for improvement in our method. Further exploration is needed to more accurately capture key visual tokens. Additionally, directly removing background information may potentially harm the model’s performance on certain tasks since the background also contains information relevant to problem-solving. In the future, we will continue researching these two areas to enhance the performance of our method further.